OpenStudy (anonymous):

A compamy will need $25,000 in7 years for a new addition. To met the goal, the company deposits money in an account today that pay 3% annula interest compund quarterly. Find the amount that should be invested to total %25,000 in 7 years?

OpenStudy (cwrw238):

25,000 = A( 1 + r/n)^(4n) where A is the amount invested, r = annual rate = 0.03, and n = number of payments per year (=4) solve for A

OpenStudy (anonymous):

??

OpenStudy (cwrw238):

this is the formula for amount obtained by compound interest

OpenStudy (cwrw238):

plug in values of r, n and solve for A can u do this?

OpenStudy (anonymous):

So what is the set up for the equation?

OpenStudy (cwrw238):

25,000 = A(1 + (0.03)/4)^28

OpenStudy (cwrw238):

A = 25,000 / (1 + (0.03)/4)^28

OpenStudy (anonymous):

Sorry, I am not getting this

OpenStudy (cwrw238):

have you seen this formula before?

OpenStudy (anonymous):

not really, I am not good with formulas

OpenStudy (cwrw238):

compound interest accumulates every so often - in this case its every 3 months (quarterly) compound means that after each period the interest on the accumulated total is added. for example 1000 invested at 5% annually after 1 year the anount becomes 1000 + 5% 0f 1000 = 1050 after 2 years ............................ 1050 * 5% of 1050 this leads to the formula i posted in that case A = amount to be invested (which we need to find ) . we need to find this using algebra

OpenStudy (cwrw238):

a simple example is 1000 = 10 A so A = 1000/10 -but in our case its not 10 its (1 + (0.03)/4)^28

OpenStudy (anonymous):

Is the answer 250?

OpenStudy (anonymous):

Ok, so the had to put 750 deposit, so after that they would have to put down 250 every year for 7 years to get 25,000????

OpenStudy (cwrw238):

no the answer is A = 25,000 / (1 + (0.03)/4)^28 which you can do on a calculator it comes $20,280.48

OpenStudy (anonymous):

So the company has to invest 20,280.48 dollars?

OpenStudy (cwrw238):

theres just one investment and this accumulates over 7 years to 25,000

OpenStudy (cwrw238):

yea - oh its 20,280.49 - my typo

OpenStudy (anonymous):

20,280.49 to the nearest dollar

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

nearest cent

OpenStudy (anonymous):

nearest dollar

OpenStudy (anonymous):

well as neeed

OpenStudy (cwrw238):

no sir - its 20,280 dollars and 49 cents

OpenStudy (cwrw238):

to nearest dollar its 20,280

OpenStudy (cwrw238):

- i don't think the companies too worried about cents...

OpenStudy (cwrw238):

hope this has helped

OpenStudy (anonymous):

20,280.00 correct?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

You have really helped me out alot, thank you for all your help

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

(;